40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 26 22:45:14 2013, manasa, Update, General, Vent progress 
    Reply  Tue Jul 30 00:21:15 2013, Koji, Update, General, Vent progress 
       Reply  Tue Jul 30 11:30:55 2013, manasa, Update, General, ITMX oplev 
          Reply  Tue Jul 30 13:41:35 2013, Koji, Update, General, ITMX oplev 
             Reply  Tue Jul 30 15:38:24 2013, manasa, Update, General, ITMX oplev 
Message ID: 8936     Entry time: Tue Jul 30 00:21:15 2013     In reply to: 8929     Reply to this: 8937
Author: Koji 
Type: Update 
Category: General 
Subject: Vent progress 

[Koji, Manasa, Sujan] 

Tomorrow we'll make final checks of the optics inside the chamber.
Then we will pump down the chamber.

- AS beam steering
The AS beam was aligned on the steering mirror (Piezo Jenna i.e. PZT4).
At this point we already found the beam on the AS CCD.
After checking the spots in the OMC chamber, we determined the beam is not clipped inside.
 
On the AS table, there is a weird thick lens with short focal length (like f=37.5mm), this actually clipped the beam significantly
[I take the blame for it from the last vent. The lens was installed thinking it would be temporary which it was not. -Manasa].
We removed this lens and put the CCD camera back so that the spot on the camera became smaller.
 
- SRM debiasing
 
First of all, the bias of the SRM was moved to zero.
Then the tower was rotated so that the beam is aligned in yaw.
After fine alignment of the SRM, the yaw slider was changed from -5.8 to 1.2.
The OSEMs on the SRM suspension was also checked. The OSEM insertions were
adjusted so that they show the half values of the open calibration
(the values can be found on the 40m wiki)
 
- Oplev check
 
SRM / ITMY / BS / PRM oplev paths in the chamber were checked.
The beams were aligned on the steering mirrors and the QPDs on the optical table.
The ITMX oplev was also checked from the windows without removing the door.
The beam is actually hitting the right positions of the steering mirror and the test mass
although there are lots of stray beams because of the BS oplev and some halo assciated with the ITMX He-Ne laser(!?).

- Checking the alignment
After we closed the light doors, we aligned the MC, the arms and the recycling mirrors.
Now the everything is aligned.
ELOG V3.1.3-